OpenStudy (anonymous):

I have 2 equations of 2 lines.I want to find point of intersection of those lines.How it could be done? Thanks in advance.

OpenStudy (anonymous):

Set the two equations equal to each other then solve them for where they equal 0

OpenStudy (anonymous):

Can you elaborate please?

OpenStudy (anonymous):

give us the two equations

OpenStudy (anonymous):

ok.wait.

OpenStudy (anonymous):

Solve both equations for the same variable, then set them equal and solve.

OpenStudy (anonymous):

Find where they intersect? o.o It should be a spot where they both meet kinda making an X.

OpenStudy (anonymous):

first equation = second equation you'll find the x they meet but first you gotta sum the 2 lines so it will be 1 equations then you do what i said

OpenStudy (anonymous):

Ok. here are the equations: x- 7y +5 = 0 and 3x + y -7 = 0

OpenStudy (amistre64):

to find the point of x and y that both lines have in common; we just make them equal each other.

OpenStudy (anonymous):

I would put everything in terms of y first and so you'd have y=-3x+7 and y= x/7+5/7 then set them equal to each other

OpenStudy (anonymous):

In terms of x I mean

OpenStudy (amistre64):

x- 7y +5 = 3x + y -7 ; now move one side to the other: x-3x -7y-y +5+7 = 0 ; combine the like terms -2x -8y +12 = 0 ; well, one trial down :) try anothe rmethod

OpenStudy (anonymous):

x = 2 y = 1

OpenStudy (anonymous):

ok. let me try. I 'll solve for y. wait please!!

OpenStudy (anonymous):

(2,1)

OpenStudy (anonymous):

-3x+7 = x/7+5/7

OpenStudy (amistre64):

Yeah, find a way to eliminate one of the variable; or to substitute a 'value' into the other equation; ill try elimination here x - 7y +5 = 0 <- multiply this by -3 3x +y -7 = 0 -3x +21y -15 = 0 3x + y -7 = 0 ; now add them together ------------------ 22y -22 = 0 ; now its easier to solve for just the 'y'

OpenStudy (anonymous):

Got it. x=2 and y=1. So, is this the intersection point? what is logic behind it?

OpenStudy (amistre64):

the logic is that in order for these lines to cross; they have to have at least one point in common

OpenStudy (amistre64):

at that point the x and y values are the same

OpenStudy (anonymous):

"at that point the x and y values are the same" for both the lines, right?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

ok. now I can see through. Thanks everybody for helping me out.I appreciate.it.

OpenStudy (amistre64):

☺ youre welcome